The longer sides of a rectangle are six inches more than twice the length of the shorter sides. The perimeter of the rectangle is 50 inches. What are the dimensions of the rectangle?
Answer by rfer(16322)

2x+2(2x+6)=50
6x+12=50
6x=50-12
x=38/6
x=6.3
2x+6=18.6